5. Identify the following: k (number of groups or samples) N, (number of cases within each group or sample) N (total number of cases) df (between-groups degrees of freedom)= How is it calculated? df (within-groups degrees of freedom) How is it calculated? critical value for F (using the F table) SSB (Sum of Squares Between groups) = ssw (Sum of Squares Within groups) Mean square between= How is it calculated? Mean square within= How is it calculated? F ratio (test statistic)= How is it calculated? 6. Is the test statistic within the critical region? YES NO 7. Can you reject the null? YES/NO 8. Is there a statistically significant difference between at least one mean and the others? YES/NO 9. What do your findings suggest is the answer to Research Question 1 (stated above)? (Be specific and mention the variables in this problem.)
Descriptive Statistics: Descriptives JobAiely 95% Conlidence Inlerval lor Mean Maximum Upper Bound N Mean Std. Deviation Std. Error Lower Bound Minimum Control 20 5.0939 7.1061 6.1000 2.14966 48068 3.00 10.00 Pop Talk 20 46396 4.1000 2.07491 3.1289 5.0711 1.00 7.00 Realism 20 6.7500 2.07428 46382 5.7792 7,7208 3,00 10.00 5.6500 5.0412 Total 60 2.35656 30423 6.2588 1.00 10.00 ANOVA Hypothesis Test Results: ANOVA JobAxiely Sum of Squares dr Mean Square Sig. Botwoon Groups 2 8.651 001 78.300 38.150 Within Groups 57 251.350 4.410 Totul 327.850 59
Answer #1

5) h=3 SSB76.3 N-60 SSW=261. 35 MSB= SSB/dl1= 38.15 MSW SSWldl2 MSB dl1= 3-1-2 4.416 60-3-57 F= MSW Citcal F = F005; 2,57ニ315